OpenStudy (anonymous):

-1 / (tanx - secx) = (1+ sinx) / cosx

OpenStudy (mertsj):

Is this an identity you are trying to prove?

OpenStudy (anonymous):

yes It is

OpenStudy (mertsj):

I would start out on the left side. Replace tan with sin/cos. Replace sec with 1/cos Add the two together and then multiply numerator and denominator by cos. See where that takes you.

OpenStudy (anonymous):

ok I got -1 over cosx/sinx - 1

OpenStudy (mertsj):

That can't be right. If you multiply numerator and denominator by cos, the numerator will be -cos since -1 times cos is -cos

OpenStudy (anonymous):

Oh yeah your right. ok its -cosx over sinx -1

OpenStudy (mertsj):

Ok. Now I think I would multiply numerator and denominator by sinx + 1

OpenStudy (anonymous):

ok I have -cosx(sinx+1) over sin^2x + 1

OpenStudy (mertsj):

Nope. (sinx-1)(sinx+1)=sin^2-1

OpenStudy (mertsj):

You should have: \[\frac{-\cos x(\sin x+1)}{\sin ^2x-1}\]

OpenStudy (anonymous):

yes your right

OpenStudy (mertsj):

Now change -cosx to cos x and change the denominator to its opposite.

OpenStudy (mertsj):

Then in the denominator you will have 1-sin^2x Replace that with cos^2x and cancel cos x

OpenStudy (anonymous):

that just leaves me sinx + 1. I need a cosx on the bottom

OpenStudy (anonymous):

Oh nevermind its squared on the bottom

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

I got it to equal!!! Thank you very much

OpenStudy (mertsj):

yw